7/3x+1/3x=4x+55/3+7/3x
Answer by rfer(16322)

8/3x=12/3x+55/3+7/3x
-19/3x+8/3x=55/3
-11/3x=55/3
x=55/3*-3/11
x=-165/33
x=-5